# Preference Curriculum: LLMs Should Always Be Pretrained on Their Preferred Data

### v1

@@ B.1 Proportion Preference Annealing Experiment
In this subsection, our goal is to systematically explore the model’s preference for sample mixing ratios based on pd values (i.e., ppl differences) at different pretraining stages. This process is crucial for understanding the dynamic changes in model preferences during the learning process and provides empirical evidence for optimizing curriculum learning strategies.
Firstly, we construct the dataset required for the annealing experiment. Based on the median of pd values, samples are divided into two groups: pd-small and pd-large. The median is set as the threshold, with samples having pd values less than the median classified as pd-small, and those greater as pd-large. From the pd-small group, we construct annealing datasets with varying concentration ratios. Setting the concentration ratio of pd-small samples as $c$, we create 11 different annealing datasets where $c$ takes values of 0%, 10%, 20%, …, 100%. To enhance data diversity, each dataset is supplemented with 30% of samples that share the same distribution as the pretraining data, and the mixed samples remain consistent across all datasets. This design ensures the robustness and comparability of the experimental results.
The annealing experiments are conducted at various stages of model pretraining to evaluate the model’s preference for sample mixing ratios at different training progressions. We perform experiments on checkpoints from 8 pretraining stages, corresponding to completion levels of 0%, 12.5%, 25%, 37.5%, 50%, 62.5%, 75%, 87.5%, and 100%. At each checkpoint, we evaluate the model using all 11 annealing datasets and record the model’s performance across different mixing ratios.
By conducting annealing experiments at checkpoints throughout the pretraining stages, we obtain a series of preference data points $(p,b)$, where $p$ represents the pretraining completion level, and $b$ denotes the model’s preference for the concentration of pd-small samples at that stage. Specifically, for each stage $p$, $b$ is defined as the concentration ratio of pd-small samples that optimizes model performance, i.e.,
$b=\arg\max_{c}\mathcal{M}(c)$ (9)
where $\mathcal{M}(c)$ denotes the comprehensive performance metric of the model on the annealing dataset with pd-small sample concentration $c$.
In our study, to fit the changes in concentration preferences of large language models during pretraining, we employ the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP) method. This approach captures the nonlinear and complex variations in preference for different pd data ratios while maintaining monotonicity and smoothness, avoiding the oscillation issues that traditional polynomial interpolation may encounter.
Given experimental data points $(p_{i},b_{i})$, describing the relationship between pretraining completion $p$ and concentration preference $b$, the PCHIP method constructs local cubic polynomials to ensure monotonicity and smoothness within each interval. Specifically, for each interval $[p_{i},p_{i+1}]$, PCHIP generates a cubic polynomial:
@@
We apply the interpolation function to a uniformly distributed set of points from 0 to 1 to obtain a continuous function curve of concentration preference $b$ as it varies with pretraining completion $p$. To ensure the physical significance of the results, we constrain the values of the fitted curve between 0 and 1. This method effectively captures the trend of model preferences for datasets at different pretraining stages, providing a reliable foundation for subsequent analysis.
@@
2:Output: Optimized preference function $f^{*}(p)$
@@
4:     Proportion Preference Annealing Experiment:
5:     Divide $\mathcal{D}$ into pd-small and pd-large based on median PD values
6:     Construct 11 annealing datasets with pd-small concentration $c\in\{0\%,10\%,\ldots,100\%\}$
7:     Supplement each dataset with 30% of samples from the original distribution
8:     for each checkpoint $p\in\{0\%,12.5\%,\ldots,100\%\}$ do
9:         Evaluate model performance $\mathcal{M}(c)$ on all annealing datasets
10:         Record preference $b=\arg\max_{c}\mathcal{M}(c)$
@@
12:     Fit Preference Function:
13:     Use PCHIP to fit preference function $b=f(p)$ from data points $(p_{i},b_{i})$
14:     Constrain $f(p)$ to $[0,1]$ for physical significance
15:     Iterative Curriculum Learning:
16:     Apply $f(p)$ to determine optimal pd-small proportion during training
@@
18:     Check for Convergence:
19:     if change in $f(p)$ between iterations $<\epsilon$ then
20:         Converged
21:     end if
22:end while
23:return $f^{*}(p)$
B.2 Iterative Curriculum Learning
After completing the proportion preference annealing experiment and successfully fitting the preference function $b=f(p)$, we apply this function to optimize curriculum learning strategies, guiding the pretraining process of the model. Specifically, based on the fitted function $f(p)$, we dynamically determine the optimal proportion $b$ of pd-small samples during training according to the current pretraining completion $p$. This approach allows us to simulate a dynamic sample mixing ratio allocation through the orderly processing of static pretraining data.
However, it is important to note that the integral of the fitted preference function over the interval $[0,1]$ may not equal 0.5. This implies that, under this configuration, the amounts of pd-small and pd-large data used may not be equal. To ensure the reasonable utilization of all data, we calculate the integral of the function over $[0,1]$ to determine the quantile threshold for dataset division. The specific formula is:
[…]
This integral value $\alpha$ guides us in adjusting the data allocation ratio.
After pretraining is completed, we conduct the proportion preference annealing experiment again to obtain updated preference data points $(p_{i},b_{i})$, and refit the preference function $f(p)$ based on these data. This process continues iteratively until the following termination condition is met: the change in the preference function between two consecutive iterations is below a predefined threshold. This method ensures that the model’s concentration preference for the dataset is precisely adjusted and optimized during pretraining.
@@ B.2 Iterative Curriculum Learning
Figure 11: Few-shot downstream performance on various benchmarks with respect to pretraining iterations for Random and (PDPC, PD, S.). We train a 3B model over 1T tokens, demonstrating superior performance with our approach.
Appendix C Experimental Details
C.1 Detailed performance on the benchmarks
Figure 11 illustrates the changes in detailed metrics as training progresses, showing few-shot downstream performance on various benchmarks with respect to pretraining iterations for Random and (PDPC, PD, S.). We train a 3B model over 1T tokens, demonstrating superior performance with our approach.
C.2 PD distribution across different domains
As illustrated in Figure 12, large size discrepancies among RMs result in broader PD distributions, which enhance data differentiation. This finding is supported by ablation tests, where the 100M-1.3B PD calculations yielded the best results. Additionally, PD maintains a consistent distribution across domains; for instance, the PD between 100M and 700M models typically follows a normal distribution with a mean of 0.27. This stability makes PD a reliable metric, as it remains unaffected by changes in data properties. Consequently, organizing training data by PD ensures it does not skew towards specific sub-domains, allowing the model to encounter diverse data at every stage.
Figure 12: PD distribution across different domains.
(a) PPL of 100M RM
(b) PPL of 700 RM
Figure 13: PPL distribution across different PD data
C.3 PPL Distribution
As shown in Figure 13, examining the perplexity distribution across data with varying PD values reveals that samples with lower PD exhibit lower perplexity. This observation aligns with the trends illustrated in Figure 5(c), where training with smaller PD leads to a rapid decrease in training loss during the initial stages, providing the model with a clear direction for gradient optimization. Notably, even when handling data with larger PD values in later stages, the model maintains a steady rate of loss reduction, ultimately achieving a minimized training loss.
C.4 Prompts for Case Study
The prompt and specific rules used in Section 3.5 to analyze the linguistic features of data across different PD intervals is as follows.
@@ C.4 Prompts for Case Study
Prompts for Property Recognition 1. Does the text contain polysemous words? Polysemous words may make understanding more difficult. 2. Does the text use specialized terminology? Specialized terminology may require specific domain knowledge to understand. 3. Does understanding the text require specific cultural background knowledge? Cultural background dependence may increase the complexity of understanding. 4. Does the text require logical reasoning to understand? Logical reasoning adds depth to understanding. 5. Does the text contain elements of humor? Humor may affect the way the text is understood. 6. Does the text explore ethical or moral issues? This may increase the depth of thought. 7. Does the text use complex sentence structures? Complex sentence structures may increase the difficulty of understanding. 8. Does the text contain scientific or technical concepts? These concepts may require specific knowledge to understand. 9. Does the text express obvious emotional tones? Emotional tones may affect the understanding of the text. 10. Does understanding the text require additional background knowledge? Background knowledge requirements may affect the comprehensibility of the text.
@@ Appendix D Case Study
Table 4: Samples are divided into 10 Probability of Default (PD) quantiles, with two samples representing each quantile.
@@
Sample 1: The need to practice good self-care doesn’t change in this working environment, but how you accomplish this goal might. Much of Arel’s own self-care regimen needed to be adjusted."I was used to weekly massage and monthly chiropractic care. That was gone," she explains. "I am used to runs and yoga and time to meditate in complete silence. That was gone, too." Although there was an onboard fitness center available, Arel often found that she couldn’t get a workout in during the day and was too tired at the end of the day to be invested in the results.Related…
Sample 2: I have to ask you, why’d you–wha–wha–why are you peeing right here?Creepy Guy: What?Kumar: I mean… why’d you pee right next to me when you could like, choose that bush, or–?Creepy Guy: Well, this bush looked like I should pee on it. Why are you peeing on it?Kumar: Well, no one was here when I chose this bush.Creepy Guy: Oh, so you get to pee on it and no one else does? Huh?Kumar: No, it’s just–I just–Creepy Guy: This your bush? You have a special bond with this bush?Kumar: No, I just thought that–Creepy Guy: You the king of the forest…
@@
Sample 1: Maybe it just sagslike a heavy load. Or does it explode?   by Langston HughesIn 1849, Elizabeth Blackwell became the first woman to graduate from a U.S. medical school in N.Y.In 1864, Rebecca Lee Crumpler became the first black woman to graduate from a U.S. medical school in New England.In 1915, women represented approximately 5% of the physician workforce in the U.S.In 1983, women represented approximately 1/3 of U.S. medical school matriculants.In 1976, Grady doctor and cardiologist Nanette K. Wenger is recognized as one of Time Magazine’s Women of the year, and in 1998 she is named Chief of Cardiology.In 2006, women represented nearly one half of all U.S. medical …
Sample 2: FILED UNDER SEAL PURSUANT TO PROTECTIVE ORDER rise to a direct infringement claim against it. See, e.g., Akami Techs., Inc. v. Limelight Networks, Inc., 797 F.3d 1020, 1023 (Fed. Cir. 2015) (noting entities are liable for performance they control). The evidence further shows that Badoo Software Limited and Badoo Limited are also intimately involved in Badoo Trading’s creation and ownership of the infringing Bumble application. Ex. A at Recital (c). Badoo Limited’s records indicate that its "principal activity" is "to provide mobile development services to Badoo Software Limited, another entity in the Worldwide Vision group and they like to…

### v2

@@ B.1 Proportion Preference Annealing Experiment
In this subsection, we aim to systematically explore the model’s preference for data mixing ratios based on PD at different pretraining stages. This process is crucial for understanding the dynamic changes in model preferences during the pretraining process and provides empirical evidence for optimizing curriculum learning strategies.
Firstly, we construct the dataset required for the annealing experiment. Based on the median of PD values, samples are divided into two parts: low-PD data and high-PD data. We create 11 different annealing datasets where the proportion of low-PD data takes values of 0%, 10%, 20%, …, 100%. To enhance data diversity, each dataset is supplemented with 30% of samples that share the same distribution as the pretraining data, and the mixed samples remain consistent across all datasets. This design ensures the robustness and comparability of the experimental results.
The annealing experiments are conducted at various stages of model pretraining to evaluate the model’s preference for data mixing ratios at different training progressions. We perform experiments on checkpoints from 8 pretraining stages, corresponding to pertraining progress of 0%, 12.5%, 25%, 37.5%, 50%, 62.5%, 75%, 87.5%, and 100%. At each checkpoint, we evaluate the model using all 11 annealing datasets and record the model’s performance across different mixing ratios.
By conducting annealing experiments at checkpoints throughout the pretraining stages, we obtain a series of preference data points $(p,b)$, where $p$ represents the pretraining progress, and $b$ denotes the model’s preference for the proportion of low-PD data at that stage. Specifically, for each stage $p$, $b$ is defined as the proportion of low-PD data that optimizes model performance, i.e.,
$b=\arg\max_{\beta}\mathcal{M}(\beta)$ (9)
where $\mathcal{M}(\beta)$ denotes the comprehensive performance metric of the model on the annealing dataset with the proportion of low-PD data $\beta$.
In our study, to fit the changes of preference of LLMs during pretraining, we employ the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP) method. Given experimental data points $(p_{i},b_{i})$, the PCHIP method constructs local cubic polynomials to ensure monotonicity and smoothness within each interval. Specifically, for each interval $[p_{i},p_{i+1}]$, PCHIP generates a cubic polynomial:
@@
We apply the interpolation function to a uniformly distributed set of points from 0 to 1 to obtain a continuous function curve of concentration preference $b$ as it varies with pretraining progress $p$. We constrain the values of the fitted curve between 0 and 1. This method effectively captures the trend of model preferences for datasets at different pretraining stages, providing a reliable foundation for subsequent analysis.
@@
2:Output: Iteratively trained model $\theta_{K}$
@@
4:     Partition $\mathcal{D}$ into $2$ sub-domains $A_{PD}^{low}$ and $A_{PD}^{high}$.
5:     Construct the annealing dataset $\mathcal{D}_{i}$ with the proportion of low-PD data set to $\beta_{i}$:
6:     $\beta_{i}\in\{0\%,10\%,\ldots,100\%\}$
7:     for $p$ in $\{0\%,12.5\%,\ldots,100\%\}$ do
8:         Retrieve the model checkpoint $\theta_{p}$
9:         Evaluate model performance $\mathcal{M}_{p}(c)$ on $\{\mathcal{D}_{i}\}_{i=0}^{10}$
10:         Record preference $b_{p}=\arg\max_{c}\mathcal{M}_{p}(c)$
@@
12:     Use PCHIP to fit $b=f(p)$ from data points $(p_{i},b_{i})$
13:     if change in $f(p)$ between iterations $<\epsilon$ then
14:         Break
15:     end if
16:     for $k=0$ to $K-1$ do
17:         Calculate pretraining progress $p=\frac{k}{K}$
18:         Get the proportions vector:
19:         $[\alpha_{1},\alpha_{2}]\leftarrow[f(p),1-f(p)]$
20:         Sample from the two domains to form $B_{k}$:
21:         $B_{k}=\{x\sim A_{PD}^{low}\}_{\alpha_{1}\cdot N}\cup\{x\sim A_{PD}^{high}\}_{\alpha_{2}\cdot N}$
22:         Train the model on $B_{k}$ and update $\theta_{k}$
23:     end for
24:end while
B.2 Iterative Curriculum Learning
After completing the proportion preference annealing experiment and successfully fitting the preference function $b=f(p)$, we apply this function to optimize curriculum learning strategies, guiding the pretraining process of the model. Specifically, based on the fitted function $f(p)$, we dynamically determine the optimal proportion $b$ of low-PD data during training according to the current pretraining completion $p$.
However, it is important to note that the integral of the fitted preference function over the interval $[0,1]$ may not equal 0.5. This implies that, under this configuration, the amounts of low-PD and high-PD data used may not be equal. To ensure the reasonable utilization of all data, we calculate the integral of the function over $[0,1]$ to determine the quantile threshold for dataset division. The specific formula is:
@@ B.2 Iterative Curriculum Learning
where $\alpha$ guides to adjust the data allocation ratio.
After pretraining is completed, we conduct the proportion preference annealing experiment again to obtain updated preference data points $(p_{i},b_{i})$, and refit the preference function $f(p)$ based on these data. This process continues iteratively until the following termination condition is met: the change in the preference function between two consecutive iterations is below a predefined threshold. This method ensures that the model’s data preference is precisely adjusted and optimized.
Appendix C Experimental Details
C.1 PPL Distribution of low-PD and high-PD data
As shown in Figure 10, examining the perplexity distribution across data with varying PD values reveals that samples with low PD exhibit lower perplexity. This observation aligns with the trends illustrated in Figure 5(c), where training with low-PD data leads to a rapid decrease in training loss during the initial stages, providing the model with a clear direction for gradient optimization. Notably, even when handling with high-PD data in later stages, the model maintains steady loss reduction, ultimately achieving a minimized training loss.
(a) PPL of 100M RM.
(b) PPL of 700 RM.
Figure 10: PPL distribution of low-PD and high-PD data.
C.2 Spearman correlation coefficient of PDs from different RM sizes
We evaluate the Spearman correlation coefficients between different PD types, as shown in Figure 11, and find a strong correlation among PDs derived from RMs of varying sizes, which indicates calculating PD with smaller RMs produces results consistent with larger RMs, optimizing computational resources.
C.3 PD distribution across different domains
As illustrated in Figure 12, large size discrepancies among RMs result in broader PD distributions, which enhance data differentiation. This finding is supported by ablation tests, where the 100M-1.3B PD calculations yielded the best results. Additionally, PD maintains a consistent distribution across domains. This stability makes PD a reliable metric. Organizing training data by PD ensures it does not skew towards specific sub-domains, allowing the model to encounter diverse data at every stage.
Figure 12: PD distribution across different domains.
C.4 Detailed performance on the benchmarks
In this section, we explore the detailed performance metrics across various benchmarks. Figure 13 illustrates how these metrics evolve during training, focusing specifically on few-shot downstream performance. We compare the performance of Random and PDPC-PD-S., across pretraining iterations.
Our experiments involved training a model with 3 billion parameters on a dataset containing 1 trillion tokens. This large-scale training setup effectively demonstrates our approach’s superior performance. The results highlight the gains in accuracy and performance achieved by our method, showcasing its clear advantages over Random. We affirm the potential of the PDPC-PD-S. methodology in enhancing model performance, particularly when faced with diverse and challenging benchmarks.
@@ C.4 Detailed performance on the benchmarks
Figure 13: Few-shot downstream performance on various benchmarks with respect to pretraining iterations for Random and PDPC-PD-S.. We train a 3B model over 1T tokens, demonstrating superior performance with our approach.
C.5 Prompts for Case Study
The prompt and specific rules used in Section 3.5 to analyze the linguistic features of data across different PD intervals are as follows.
@@ C.5 Prompts for Case Study
{Property} is from one of the following rules:
Rules for Property Recognition 1. Does the text contain polysemous words? Polysemous words may make understanding more difficult. 2. Does the text use specialized terminology? Specialized terminology may require specific domain knowledge to understand. 3. Does understanding the text require specific cultural background knowledge? Cultural background dependence may increase the complexity of understanding. 4. Does the text require logical reasoning to understand? Logical reasoning adds depth to understanding. 5. Does the text contain elements of humor? Humor may affect the way the text is understood. 6. Does the text explore ethical or moral issues? This may increase the depth of thought. 7. Does the text use complex sentence structures? Complex sentence structures may increase the difficulty of understanding. 8. Does the text contain scientific or technical concepts? These concepts may require specific knowledge to understand. 9. Does the text express obvious emotional tones? Emotional tones may affect the understanding of the text. 10. Does understanding the text require additional background knowledge? Background knowledge requirements may affect the comprehensibility of the text.
@@ Appendix D Case Study
Table 4: Samples are divided into 10 PD quantiles, with two samples representing each quantile.
@@
Sample 1: The need to practice good self-care doesn’t change in this working environment, but how you accomplish this goal might. Much of Arel’s own self-care regimen needed to be adjusted."I was used to weekly massage and monthly chiropractic care. That was gone," she explains. "I am used to runs and yoga and time to meditate in complete silence. That was gone, too." …
Sample 2: I have to ask you, why’d you–wha–wha–why are you peeing right here?Creepy Guy: What?Kumar: I mean… why’d you pee right next to me when you could like, choose that bush, or–?Creepy Guy: Well, this bush looked like I should pee on it. Why are you peeing on it?Kumar: Well, no one was here when I chose this bush.Creepy Guy: Oh, so you get to pee on it and no one else does? Huh?Kumar: …
@@
Sample 1: Maybe it just sagslike a heavy load. Or does it explode?   by Langston HughesIn 1849, Elizabeth Blackwell became the first woman to graduate from a U.S. medical school in N.Y.In 1864, Rebecca Lee Crumpler became the first black woman to graduate from a U.S. medical school in New England.In 1915, women represented approximately 5% of the physician workforce in the U.S.In 1983, women represented approximately 1/3 of U.S. medical school matriculants …
Sample 2: FILED UNDER SEAL PURSUANT TO PROTECTIVE ORDER rise to a direct infringement claim against it. See, e.g., Akami Techs., Inc. v. Limelight Networks, Inc., 797 F.3d 1020, 1023 (Fed. Cir. 2015) (noting entities are liable for performance they control). The evidence further shows that Badoo Software Limited and Badoo Limited are also intimately involved in Badoo Trading’s creation and ownership of the infringing Bumble application…